OpenStudy (anonymous):

Help please!! Kersha has two jobs. During the day she works as an office clerk, and in the evening she works as a cashier. Her office job pays her $12.00 per hour. Her cashier job pays her $8.25 per hour. In one week Kersha worked 55 hours. She earned a total of $585. How many hours did Kersha work in each job?

OpenStudy (anonymous):

alright to find this type of problem you have to have a formula. pick two variables for the two jobs to represent the hours worked in each/

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me know what you pick.

OpenStudy (anonymous):

Well this was my answer. i wasnt sure if i was correct. The correct equation is 12c + 8.25(55 - c) = 585

OpenStudy (anonymous):

gee im dumb.this was foranother problem maybe thats why i wouldnt get it

OpenStudy (anonymous):

well what i meant to say was 35 hours clerk, 20 hours cashier.

OpenStudy (anonymous):

Okay, look: your person works x hours in job one and z hours in job 2. job 1 earns $12/h. job two earns $8.50/h. correct? and she worked for 55 hours total. and earned $585

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, if you want to check if thats the right answer, play Plug N Chug. Your equation is 12x + 8.50z = 585 so plug in 35 for x and 20 for z. I hope this helped, i gotta go now.

OpenStudy (anonymous):

alright thanks!

OpenStudy (wolf1728):

you'll need a second equation to avoid a trial and error solution. Here is your second equation: x + z = 55